OpenStudy (anonymous):

a= -2 b= -3 c= 2 x= 2.1 y= 3 and z= -4.2 I y - 2z I -3

OpenStudy (calculusfunctions):

Substitute y = 3 and z = -4.2 into the expression and tell me what you get.

OpenStudy (anonymous):

So i asked for help and you tell me to do it myself?

OpenStudy (calculusfunctions):

No sorry, you're right. Let me start over.

OpenStudy (calculusfunctions):

Substitute y = 3 and z = -4.2 into the expression to obtain |3 − 2(-4.2)| − 3 = |3 + 8.4| − 3 = |11.4| − 3 = 11.4 − 3 = 8.4 Was that helpful?

OpenStudy (calculusfunctions):

If you need further clarification on any step, please don't hesitate to ask.

OpenStudy (anonymous):

thanks!(:

OpenStudy (calculusfunctions):

Anytime!